# FireEdit: Fine-grained Instruction-based Image Editing via Region-aware Vision Language Model

### v1

@@ 1 Introduction
As a critical area of research in multimedia interaction, image editing has consistently garnered significant attention from both academia and industry. The emergence of diffusion models [16, 15, 47] and autoregressive models [49, 50] has reshaped the fiels of image generation [19, 33, 7, 32] and video generation [23, 28, 55, 26, 56]. This advancement provides enhanced control and flexibility for generative image editing, enabling the development of numerous commercial products. By transferring the discrepancies between input and target text descriptions to a cross-modal implicit mapping of intermediate latent features, the desired regions of real images can be modified [14, 39, 24, 3, 34]. Precise editing is achievable when region masks are available; however, fine-grained text descriptions and region masks are often difficult to obtain.
@@
Instruction-based image editing methods circumvent these seemingly essential conditions and are highly sought after by the community for their simplicity, user-friendliness, and efficiency. IP2P [4] is a pioneering work in instruction-based editing, allowing users to express their modification intentions for a specified image using straightforward natural language. Recent methods strive to generate high-quality input-target-instruction triples at scale [61, 48, 22, 64, 12] and leverage the remarkable reasoning and comprehension capabilities of LLMs [11, 21, 60]. Despite the significant progress these methods have made in advancing instruction-based image editing, there are still shortcomings in the following scenarios: 1) complex instructions; 2) intricate scenes (e.g., multi-object environments and urban streets); and 3) fidelity degradation, which includes undesirable changes in style, texture, color, and other visual details. As illustrated in Figure 1, current methods either exhibit inaccurate localization or inadvertently alter unintended objects. These challenges stem from two primary factors. First, the CLIP text encoder [45] utilized by Stable Diffusion [47] struggles to extract the desired modifications from the instructions provided by general users. In complex scenarios, even when users invest considerable time crafting their instructions, aligning the text with the editing regions in the input images remains challenging. Second, the data pairs used for training are predominantly synthesized by text-to-image (T2I) generation models [47, 14, 44, 9], which inevitably introduces implicit biases [27, 10]. Consequently, models trained on such ambiguous or imprecise data often struggle to accurately locate the editing object and lose high-frequency details in unintended regions of the input image.
@@
To address these limitations, we propose FireEdit, a novel end-to-end, instruction-based image editing method. FireEdit primarily comprises a vision language Model (VLM) and a diffusion model. The VLM converts user-provided editing instructions into an edit representation that can be comprehended by the generative model. Previous methods [11, 21] have employed VLM to improve the model’s understanding of editing instructions. MGIE [11] utilizes VLM to generate concise expression instructions, enhancing simple instructions with greater visual imagination. SmartEdit [21] addresses complex instructions by implementing a bidirectional interaction module that promotes mutual understanding between visual features and VLM outputs. Although these methods have achieved commendable results, they frequently encounter a trade-off between adherence to instructions and semantic consistency. Unlike the standard practice of adopting image tokens alongside text tokens as input to Vision-Language Models (VLMs), we provide fine-grained visual features to ground textual instructions, thereby enhancing the reasoning capabilities of the VLM. Our insight is that directly utilizing Large Language Models (LLMs) to interpret editing intentions is not sufficiently effective. Consequently, we decouple the process of understanding instructions into two distinct tasks: locating the region of interest and associating the relevant text. We observe that the perceptual skills required for localization are not the strong suit of LLMs. By integrating fine-grained region features into the tokenization process of image features, we can complement the capabilities of LLMs. Specifically, we employ an open-vocabulary object detector [66] to identify potential regions of interest and tokenize each region into region tokens, which serve as a supplement to standard image tokens. The enhanced visual tokens, along with instruction tokens, form a new input for the LLM, which implicitly locates the editing regions and articulates the modification intentions through the LLM’s powerful reasoning capabilities.
@@
Once we obtain the editing representation from the LLM response, a direct approach is to utilize it for interaction with the cross-modal attention maps of the diffusion model. Once we obtain the editing representation from the LLM response, a direct approach is to utilize it for interaction with the cross-modal attention maps of the diffusion model. However, due to the semantic gap between the output of VLM and the diffusion model, this approach may result in insufficient guidance. To address this issue, we introduce a Time-Aware Target Injection (TATI) module. As previously noted [63, 18], the diffusion model emphasizes low-frequency structures (e.g., edges and layouts) in the early stages of the denoising process, while it focuses on high-frequency details (e.g., textures and colors) in the later stages. Therefore, we propose TATI to associate the desired editing region and object shape at the low-frequency semantic level. Subsequently, during the later stages of denoising, TATI captures high-frequency details, linking the attributes of the target. To maintain spatial detail consistency between the source image and the edited image, we propose the Hybrid Visual Cross Attention (HVCA) module. HVCA integrates multi-scale visual features to deliver rich visual details, which may play a crucial role in preserving the content of non-edited areas during the intermediate feature evolution process in diffusion models. We utilize learnable queries to filter out the detailed features that need to be preserved through interaction with the text features.
@@ 2.1 Instruction-based Image Editing
The instruction-based image editing task involves using human natural language to modify specified content in the input image while minimizing the distortion of other undesired areas. Compared to the text-to-image generative task, the editing task typically entails targeted alterations to specific entities or characteristics within the image. Recently, there are some excellent works [2, 4, 61, 8, 30, 14, 52, 31] on image editing base on diffusion model. For instance, Prompt-to-Prompt [14] adjusts the cross-attention features according to the difference between the source caption and the target caption. Building on this approach, IP2P [4] fine-tunes stable diffusion on a large-scale dataset of source-target-instruction triplets constructed by Prompt-to-Prompt [14], enabling the model to perform editing under the guidance of simple natural instructions. Subsequently, MagicBrush [61] further improves the model’s ability for editing with a real-world dataset. HQ-Edit[22] utilizes DALL-E3 and GPT-4V [41] to construct a high-quality synthetic dataset for fine-tuning IP2P. InstructDiffusion [13] extends this instruction editing paradigm, training a general model capable of handling various visual tasks across multiple datasets. UltraEdit [64] scales data by combining SDXL with advanced inpainting methods [40, 2] and trains on a more powerful generative diffusion model [9]. To enhance the editing capabilities for complex instructions, SmartEdit [21] and MGIE [11] introduce VLMs to assist in training diffusion models, facilitating understanding and reasoning for complex instructions. Our method aims to achieve precise localization of editing targets in complex scenarios while preserving high-frequency details in non-edited regions.
@@ 2.2 VLMs with Diffusion Models
Large language models (LLMs) have demonstrated impressive generalization capabilities in text generation [6, 51]. Providing visual inputs to assist text generation has shown outstanding performance in a range of downstream language-to-vision tasks [54, 29, 36, 57, 46, 37, 20]. Bridging the gap between LLMs and visual inputs is a key focus of the latest Vision-Language Models (VLMs). LLaVA [36] fine-tunes LLMs on an instruction-following dataset constructed by GPT-4. xGen-MM [57] replaces the Q-Former in BLIP-2 with a scalable visual token sampler, training LLMs on large-scale, high-quality, and diverse datasets to advance various vision-language tasks. Leveraging the strong reasoning capabilities of vision-language models, researchers have begun to introduce them into the field of image generation. GILL [25] is a pioneering work in this attempt, guiding diffusion models to generate images consistent with text by using text representations inferred from vision-language models. MGIE [11] extends this to image editing, leveraging VLMs to learn concise expression instructions. SmartEdit [21] introduces a bidirectional interactive module to facilitate mutual understanding between the text representations output by VLMs and the input image features, addressing the issue of CLIP’s text encoder’s insufficient understanding of complex instructions. ILLUM [53] establishes a unified VLM with the capability of understanding images and charts, as well as image generation and editing. Unlike previous methods that train a customized VLM, PromptFix [60] utilizes the extensive world knowledge and understanding capabilities of the existing LLaVA to enhance text prompts.
Our approach also focuses on training a dedicated VLM for editing tasks. Our insight is that a general VLM often struggles to effectively understand overly simple or complex instructions, necessitating the decoupling of the instruction understanding process into locating and associating target objects. We integrate fine-grained region tokens into standard image tokenization, providing the LLM with more visual imagination possibilities, thereby obtaining more robust and accurate text representations of user editing intentions.
@@ 4.1 Region-aware Mixed-Modal Encoding
$\mathcal{P}=\mathcal{R}(\mathcal{D}(x)).$ (3)
@@
Due to the semantic gap between the text embedding space of the diffusion model and the output of the LLM, an additional Q-Former module is required to bridge them. Finally, the edited representation $e$ obtained from the VLM will be used to guide the diffusion model in the subsequent steps.
@@
here $\Delta\theta$ is the parameter of LoRA.
@@ 4.2 Hybrid Visual Cross Attention
To prevent the loss of detail information in non-target areas, we introduce a Hybrid Visual Cross Attention (HVCA) module. As shown in Figure 3, we cascade the visual features extracted by the CLIP image encoder [45] and DINOv2 [42] as visual inputs. The use of different image encoders is intended to capture rich visual details at various pixel levels. These features are then enhanced through interaction with text-related features to obtain reinforced visual features. Our HVCA consists of $L$ attention blocks, each comprising two cross-attention layers and a feed-forward network. The IP-Adapter [58] directly injects features extracted by the image encoder into the denoising network, which makes it challenging to maintain high-frequency details, especially in the field of image editing. By introducing visual content with context-aware capabilities, we ensure that the generated images have higher fidelity while maintaining editability.
Following the IP-Adapter [58], we decouple the cross-attention layers to inject the refined visual embeddings $v$ and edited representation $e$ into the cross-attention layers of the denoising network. Specifically, given the query features $Z$, edited representation $e$, and visual embeddings $v$, the output $Z^{\prime}$ of the cross-attention can be defined as follows:
@@
where $Q=ZW_{q}$, $K_{1}=eW_{k_{1}}$, $V_{1}=eW_{v_{1}}$, and $K_{2}=vW_{k_{2}}$, $V_{2}=vW_{v_{2}}$ are query, key, and value matrices. $\lambda$ is weight factor, and $d$ represents the feature dimension.
@@ 4.4 Optimization Objectives
During the diffusion process, we follow the design principles of IP2P [4]. We concatenate the encoded image latent $\mathcal{E}(x)$ and the noisy latent $z_{t}$ as input to the denoising network $\epsilon_{\delta}$. The difference lies in that we substitute the text embeddings obtained by the CLIP text encoder with $e$, and use it along with the visual feature $v$ through the decoupled cross-attention layers of the UNet to participate in the denoising process. The optimization process for $\epsilon_{\delta}$ can be expressed as follows:
$\mathcal{L}_{diff}=\mathbb{E}_{\mathcal{E}(y),\mathcal{E}(x),c,\epsilon,t}||\epsilon-\epsilon_{\delta}(t,\mathcal{C}(z_{t},\mathcal{E}(x)),e,v)||_{2}^{2},$ (7)
where $\epsilon\sim\mathcal{U}(0,1)$, $t$ is the sampling time, $y$ is the target edited images and $\mathcal{C}$ denotes the operation of concatenation along the channel dimension. The combined total loss functions for the VLM and the diffusion model are:

### v2

@@ 1 Introduction
As a critical area of research in multimedia interaction, image editing has consistently garnered significant attention from both academia and industry. The emergence of diffusion models [16, 15, 48] and autoregressive models [50, 51] has reshaped the fields of image generation [19, 33, 7, 32] and video generation [23, 28, 56, 26, 57]. This advancement provides enhanced control and flexibility for generative image editing, enabling the development of numerous commercial products. By transferring the discrepancies between input and target text descriptions to a cross-modal implicit mapping of intermediate latent features, the desired regions of real images can be modified [14, 40, 24, 3, 35, 34]. Precise editing is achievable when region masks are available. However, fine-grained text descriptions and region masks are often difficult to obtain.
@@
Instruction-based image editing methods eliminate the reliance on these seemingly essential conditions and are widely adopted in the community for their simplicity, user-friendliness, and efficiency. IP2P [4] is a pioneering work in instruction-based editing, allowing users to express their modification intentions for a specified image using straightforward natural language. To achieve precise and high-fidelity image editing, recent methods focus on large-scale generation of high-quality input-target-instruction triples  [63, 49, 22, 66, 12, 61] and exploit the remarkable reasoning and comprehension capabilities of LLMs [11, 21, 62]. Despite the significant progress made by these methods in advancing instruction-based image editing, several shortcomings persist in the following scenarios: 1) complex instructions; 2) intricate scenes (e.g., multi-object environments and urban streets); and 3) fidelity degradation, which encompasses undesirable changes in style, texture, color, and other visual details. As illustrated in Figure 1, current methods either demonstrate inaccurate localization or inadvertently alter unintended objects. These challenges arise from two primary factors. First, the CLIP text encoder [46] employed by Stable Diffusion [48] struggles to extract the desired modifications from the instructions provided by general users. In complex scenarios, even when users invest considerable time crafting their instructions, aligning the text with the editing regions in the input images remains a challenge. Second, the data pairs used for training are predominantly synthesized by text-to-image (T2I) generation models [48, 14, 45, 9], which inevitably introduce implicit biases [27, 10]. Consequently, models trained on ambiguous or imprecise data often struggle to accurately identify the editing object and may lose high-frequency details in unintended areas of the input image.
@@
To address these limitations, we propose FireEdit, a novel end-to-end, instruction-based image editing method. FireEdit primarily comprises a Vision Language Model (VLM) and a diffusion model. The VLM converts user-provided editing instructions into an edit representation that can be comprehended by the generative model. Previous methods [11, 21] have employed VLM to improve the model’s understanding of editing instructions. MGIE [11] utilizes VLM to generate concise expression instructions, enhancing simple instructions with greater visual imagination. SmartEdit [21] addresses complex instructions by implementing a bidirectional interaction module that promotes mutual understanding between visual features and VLM outputs. Although these methods have achieved commendable results, they frequently encounter a trade-off between adherence to instructions and semantic consistency. Unlike the standard practice of adopting image tokens alongside text tokens as input to VLMs, we provide fine-grained visual features to ground textual instructions, thereby enhancing the reasoning capabilities of the VLM. Our insight is that directly utilizing LLMs to interpret editing intentions is not sufficiently effective. Consequently, we decouple the process of understanding instructions into two distinct tasks: locating the region of interest and associating the relevant text. We observe that the perceptual skills required for localization are not the strong suit of LLMs. By integrating fine-grained region features into the tokenization process of image features, we can complement the capabilities of LLMs. Specifically, we employ an open-vocabulary object detector [68] to identify potential regions of interest and tokenize each region into region tokens, which serve as a supplement to standard image tokens. The enhanced visual tokens, along with instruction tokens, form a new input for the LLM, which implicitly locates the editing regions and articulates the modification intentions through the LLM’s powerful reasoning capabilities. Introducing region tokens improves the LLM’s ability to perceive the visual world.
@@
Once the editing representation is obtained from the LLM response, a direct approach is to employ it for interaction with the diffusion model. However, due to the semantic gap between the output of VLM and the diffusion model, this approach may result in insufficient guidance. To address this issue, we introduce a Time-Aware Target Injection (TATI) module. As previously noted [65, 18], the diffusion model emphasizes low-frequency structures (e.g., edges and layouts) in the early stages of the denoising process, while it focuses on high-frequency details (e.g., textures and colors) in the later stages. Therefore, we propose TATI to associate the desired editing region and object shape at the low-frequency semantic level. Subsequently, during the later stages of denoising, TATI captures high-frequency details, linking the attributes of the target. To maintain spatial detail consistency between the source image and the edited image, we propose the Hybrid Visual Cross Attention (HVCA) module. HVCA integrates multi-scale visual features to deliver rich visual details, which may play a crucial role in preserving the content of non-edited areas during the intermediate feature evolution process in diffusion models. We utilize learnable queries to filter out the detailed features that need to be preserved through interaction with the text features. These two modules complement each other and enhance the controllability of the denoising process.
@@ 2.1 Instruction-based Image Editing
The instruction-based image editing task involves using human natural language to modify specified content in the input image while minimizing the distortion of undesired areas. Compared to the text-to-image generative task, the editing task typically entails targeted alterations to specific entities or characteristics within the image. Recently, there are some excellent works [2, 4, 63, 8, 30, 14, 53, 31] on image editing based on diffusion model. For instance, Prompt-to-Prompt [14] adjust the cross-attention features according to the difference between the source caption and the target caption. Building on this approach, IP2P [4] fine-tuned stable diffusion on a large-scale dataset of source-target-instruction triplets constructed by Prompt-to-Prompt [14], enabling the model to perform editing under the guidance of simple natural instructions. Subsequently, MagicBrush [63] further improves the model’s editing capabilities with a real-world dataset. HQ-Edit [22] utilize DALL-E3 and GPT-4V [42] to construct a high-quality synthetic dataset for fine-tuning IP2P. InstructDiffusion [13] extends this instruction editing paradigm, trains a general model capable of handling various visual tasks across multiple datasets. UltraEdit [66] scales up data by combining SDXL with advanced inpainting methods [41, 2] and trains on a more powerful generative diffusion model [9]. To enhance the editing capabilities for complex instructions, SmartEdit [21] and MGIE [11] introduce VLMs to assist in training diffusion models, facilitate understanding and reasoning for complex instructions. Our method aims to achieve precise localization of editing targets in complex scenarios while preserves high-frequency details in non-edited regions.
@@ 2.2 VLMs with Diffusion Models
Large language models (LLMs) have demonstrated impressive generalization capability in text generation [6, 52]. Providing visual inputs to assist text generation has shown outstanding performances in a range of downstream language-to-vision tasks [55, 29, 37, 58, 47, 38, 20]. Bridging the gap between LLMs and visual inputs is a key focus of the latest Vision-Language Models (VLMs). LLaVA [37] fine-tune LLMs on an instruction-following dataset constructed by GPT-4. xGen-MM [58] replaces the Q-Former in BLIP-2 with a scalable visual token sampler, trains LLMs on large-scale, high-quality, and diverse datasets to advance various vision-language tasks. Leveraging the strong reasoning capabilities of vision-language models, researchers have begun to introduce them into the field of image generation. GILL [25] is a pioneering work in this attempt, guides diffusion models to generate images consistent with text by using text representations inferred from vision-language models. MGIE [11] extends this to image editing, leveraging VLMs to learn concise expressive instructions. SmartEdit [21] introduces a bidirectional interactive module to facilitate mutual understanding between the text representations outputted by VLMs and the input image features, addressing the issue of CLIP’s text encoder’s insufficient understanding of complex instructions. ILLUM [54] establishes a unified VLM with the capability of understanding images and charts, as well as image generation and editing. Unlike previous methods that train a custom VLM, PromptFix [62] utilizes the extensive world knowledge and understanding capabilities of the existing LLaVA to enhance text prompts. Our approach also focuses on training a dedicated VLM for editing tasks. Our insight is that a general VLM often struggles to effectively understand overly simple or complex instructions, necessitating the decoupling of the instruction understanding process into locating and associating target objects. We integrate fine-grained region tokens into standard image tokenization, providing the LLM with more visual imagination possibilities, thereby obtaining more robust and accurate text representations of user editing intentions.
@@ 4.1 Region-aware Mixed-Modal Encoding
$\mathcal{P}=\mathcal{R}(\mathcal{D}(x)),$ (3)
@@
Due to the semantic gap between the text embedding space of the diffusion model and the output of the LLM, an additional Q-Former module is required to bridge them. Finally, we obtain the edited representation $e_{t}$, which is derived from $e$ through the Q-Former, is used to guide the diffusion model in the subsequent steps.
@@
where $\Delta\theta$ denotes the parameter of LoRA.
@@ 4.2 Hybrid Visual Cross Attention
To prevent the loss of viusal details in non-target areas, we introduce a Hybrid Visual Cross Attention (HVCA) module. As shown in Figure 3, we cascade the visual features extracted by the CLIP image encoder [46] and DINOv2 [43] as visual inputs. The use of different image encoders is intended to capture rich visual details at various pixel levels. These features are then enhanced through interaction with text-related features to obtain reinforced visual features. Our HVCA consists of $L$ attention blocks, each comprising two cross-attention layers and a feed-forward network. The IP-Adapter [59] directly injects features extracted by the image encoder into the denoising network, which makes it challenging to maintain high-frequency details, especially in the field of image editing. By introducing visual content with context-aware capabilities, we ensure that the generated images have higher fidelity while maintaining editability.
Following the IP-Adapter [59], we decouple the cross-attention layers to inject the refined visual embeddings $v$ and the edited representation $e_{t}$ into the cross-attention layers of the denoising network. Specifically, given the query features $Z$, the edited representation $e_{t}$, and the visual embeddings $v$, the output $Z^{\prime}$ of the cross-attention can be defined as follows:
@@
where $Q=ZW_{q}$, $K_{1}=e_{t}W_{k_{1}}$, $V_{1}=e_{t}W_{v_{1}}$, and $K_{2}=vW_{k_{2}}$, $V_{2}=vW_{v_{2}}$ are query, key, and value matrices. $\lambda$ is weight factor, and $d$ represents the feature dimension.
@@ 4.4 Optimization Objectives
During the diffusion process, we follow the design principles of IP2P [4]. We concatenate the encoded image latent $\mathcal{E}(x)$ and the noisy latent $z_{t}$ as input to the denoising network $\epsilon_{\delta}$. The difference lies in that we substitute the text embeddings obtained by the CLIP text encoder with $e_{t}$, and use it along with the visual feature $v$ through the decoupled cross-attention layers of the UNet to participate in the denoising process. The optimization process for $\epsilon_{\delta}$ can be expressed as follows:
$\mathcal{L}_{diff}=\mathbb{E}_{\mathcal{E}(y),\mathcal{E}(x),c,\epsilon,t}||\epsilon-\epsilon_{\delta}(t,\mathcal{C}(z_{t},\mathcal{E}(x)),e_{t},v)||_{2}^{2},$ (7)
where $\epsilon\sim\mathcal{U}(0,1)$, $t$ is the sampling time, $y$ is the target edited images and $\mathcal{C}$ denotes the operation of concatenation along the channel dimension. The combined total loss function for the VLM and the diffusion model is defined as